100 g fairly concentrated solution of CuSO is
electrolysed using 0.1 F electricity. The weight of resulting solution is (Atomic weight of Cu = 63.5)
  • 96.02
  • 98.6
  • 94.5
  • 91.2

Grade:12

1 Answers

Arun
4 years ago
Dear student
 
Faraday current deposit = 63.5/2 = 31.75 gm of copper
 
0.1 faraday current deposit 31.75*0.1 = 3.175 gm of copper
 
solution weight initial = 100gm
hence after electrolysis = 100- 3.175 = 96.825
 
Hope it helps